A student placed a drop of olive oil of volume 3.8*10 raise to power -2 m raise to power 3 on the surface of a still water and the drop spreads to form a circular film of radius 0.67m. Calculate the thickness of the oil film
1
Expert's answer
2019-09-30T10:08:36-0400
"V=Ah=\pi r^2h"


"h=\frac{V}{\pi r^2}=\frac{3.8\cdot 10^{-2}}{\pi (0.67)^2}=0.027\ m"